OpenStudy (anonymous):

Which equation represents the line that passes through the point (1,5) and -2?

OpenStudy (luigi0210):

Use the point-slope formula \(\Large y-y_{1}=m(x-x_{1})\)

OpenStudy (anonymous):

Give me a answer

OpenStudy (luigi0210):

That's not what I'm for.. but if you don't want help, then fine by me~

OpenStudy (anonymous):

I do shan't your help..is that I'm really lost

OpenStudy (anonymous):

Whant*

OpenStudy (luigi0210):

Well here, I'm make it simpler \(\Large (1, 5)--->(x_{1}, y_{1})\) And slope=m

OpenStudy (luigi0210):

The y and x in the equation are just variables.. so plug them in to get: \[\Large y-5=-2(x-1)\] Now just distribute and get it into slope intercept form \(\Large y=mx+b\)